OpenStudy (anonymous):

Simplify the expression: -10c - (6-3c)= ?

Nnesha (nnesha):

first distribute bracket by negative sign suppose there is one with negative sign so -1 time 6 = ?? - times -3c = ???

OpenStudy (anonymous):

-10c-6+3c=-7c-6

OpenStudy (anonymous):

-14c+-14

Nnesha (nnesha):



OpenStudy (anonymous):

Right @Princer_Jones?

OpenStudy (anonymous):

No -(+)=+(-)=- and -(-)=+

OpenStudy (anonymous):

You can prove this in Abstract algebra,

OpenStudy (uri):

-10c - (6-3c)= ? -10c-6+3c -7c-6

Nnesha (nnesha):

-1 times 6 = -6 and -1 time -3c = 3c so -10c - 6 +3c now combine like terms

OpenStudy (uri):

negative +negative=positive.

OpenStudy (uri):

negative+ positive =negative

OpenStudy (anonymous):

-7c+-6

Nnesha (nnesha):

yes so negative and positive = negative then -7c -6